# Talent Hoarding in Organizations

### v2

@@ Appendix C Theoretical Appendix - For Online Publication
Appendix D Data Appendix - For Online Publication
This section provides additional information on the firm records and survey data.
D.1 Dataset Construction
This paper uses a new dataset that consists of personnel records and application data from a large manufacturing firm. To comply with the firm’s data protection regulations, the data was assembled and anonymized in a secure data environment at the firm. Data from three data sources was combined: personnel records, talent management data, and job application data. First, following variables were collected from the firm’s personnel records: employment status, date of birth, gender, marital status, family status, parental leave status, date of entry into firm, highest educational degree attained, citizenship, function, division, location, working hours, number of direct reports, reporting distance to the CEO, managerial autonomy, base pay, bonus payments, absenteeism, date of exit from firm, date of retirement, id of manager, id of team. To create a quarterly panel dataset, the data were collapsed to the employee-by-quarter level, keeping the first observation of an employee in a given quarter. Second, following variables were collected from the firm’s talent management data: performance ratings, potential ratings, nominations to talent lists, worker assignments to trainings. Third, following variables were collected from the firm’s application data: application status and date, the outcome of an application (e.g., rejection, interview, offer), id of vacancy, id of hiring manager, hierarchy level of vacancy, division of vacancy, functional area of vacancy, location of vacancy. The three data sources were linked using a unique employee id within the firm’s secure data environment.
@@ D.2 Employee Survey
All employees in my sample were invited via e-mail by the firm’s HR department and were asked to provide their perspectives on the internal labor market at the firm. The survey received over 15,000 responses, yielding a 50% response rate. Respondents are similar to employees in my analysis sample in terms of demographics and position characteristics (Appendix Table D1). Employees described challenges regarding their internal career progression both in multiple-choice answers and in the form of free-text responses. The median response time was 13 minutes. For my analysis, I only keep respondents who took at least five minutes to complete the survey. For questions including free-text responses, I exclude the minority of respondents who did not provide any answer when presenting the share of respondents who gave a particular response. The translation of the relevant questions for this study is presented in abbreviated format: A. Please rate following six statements. “Actively applying for positions at [Company] …” {I strongly agree, I agree, Undecided, I do not agree, I totally do not agree} A.1 “… increases future promotion chances.” A.2 “… does not matter since jobs are only posted proforma.” A.3 “… would cause negative consequences by my current supervisor.” A.4 “… is seen as disloyal to my current team.” A.5 “… is appropriate once employees are unsatisfied with their job.” A.6 “… should only be done after checking in with one’s direct supervisor.” B. Which job characteristics are most important to you? Please select the two most important characteristics from the following list. {Potential for training, Potential for promotion, Pay, Flexible hours, Location, Meaningful tasks, Familiar tasks, Challenging tasks, Good relationship with colleagues, Good relationship with supervisor} C. Did someone recommend an open position to you in the past 12 months? C.1 Yes C.2 No D. You indicated that someone recommended one or more open positions to you. Who recommended these positions to you? Please select all that apply. D.1 Colleagues within your team D.2 Colleagues outside of your team D.3 Your supervisor D.4 The supervisor responsible for the open position D.5 Your HR responsible D.6 Family or friends E. At the end of this survey, we are interested in your personal opinion about current challenges and potential improvements with respect to careers at [Company]. E.1 What were the reasons why you decided in the past not to apply for internal job openings at [Company]? {free-text response} E.2 What are the main challenges that you have encountered in your career development at [Company]? {free-text response} E.3 What are some of the ways that [Company] could be helpful to you as you are planning your career? {free-text response}
@@ D.3 Manager Survey
In order to provide direct evidence on managers’ incentives and potential talent hoarding actions, a second survey was sent to managers in my sample. The invitation procedure was equivalent to the employee survey with the firm’s HR department inviting managers to participate in the survey to provide their perspectives on being a leader at the firm. The survey received a response rate of 62%, yielding over 3000 responses. Respondents are similar to managers in my analysis sample in terms of demographics and position characteristics (Appendix Table D2). The median response time was 14 minutes. For my analysis, I only keep respondents who took at least five minutes to complete the survey. The translation of the relevant questions for this study is presented in abbreviated format: A. Where do you think leaders can create the most impact for [Company]? Please rate the impact that leaders can create. {Very high impact, high impact, low impact, very low impact} A.1 Strategic orientation A.2 Vision A.3 Operational efficiency A.4 Talent development A.5 Culture and morale A.6 Innovation B. When it comes to leaders’ compensation and promotion prospects, how important do you think is a record of success in each of the following? Please rate the importance. {Very important, important, somewhat important, not important} B.1 Strategic orientation B.2 Vision B.3 Operational efficiency B.4 Talent development B.5 Culture and morale B.6 Innovation C. What are the most effective ways that leaders can support their employees’ career development? Please select all that apply. C.1 Training: Offer and recommend relevant trainings that can help employees enhance their skills and knowledge C.2 Open dialogue: Establish a regular dialogue about career goals C.3 Career planning: Introduce a career planning process that provides clear guidance for employees to take next steps in their career C.4 Visibility: Ensure that decision-makers outside the team are aware of talented employees C.5 Encouragement: Encourage employees to pursue new job opportunities throughout the organization C.6 Information about job opportunities: Notify employees of job vacancies in other areas of the organization D. What are the reasons that may prevent leaders from investing time and effort towards their employees’ career development? Please select all that apply. D.1 Leaders’ lack of knowledge about career development opportunities, trainings etc. D.2 Limited resources to invest in employee development D.3 Risk of losing talent, because developing employees makes them more attractive to others D.4 Need to prioritize short-term targets over long-term employee development D.5 Lack of interest on the part of employees E. Under which circumstances would leaders be more likely to support employees’ career development? Please select all that apply. E.1 If the employee would develop within the team E.2 If replacing employees who leave the team would be easier E.3 If leaders who develop talent were to receive more visibility and recognition E.4 If leaders who develop talent were more likely to be promoted E.5 If leaders who develop talent would receive financial incentives F. Please rate your agreement with the following statements. {I strongly agree, I agree, I do not agree, I totally do not agree} F.1 “A leader’s direct intervention (e.g., encouragement, increasing visibility) has a large impact on employees’ career development.” F.2 “Leaders’ impact on employees’ long-term career development is valued at least as much as their impact on team performance.” F.3 “Talent development entails a conflict of interest for leaders, because more developed employees are more likely to leave the team.” G. How often might leaders find themselves in situations where they need to dissuade a team member from exploring opportunities in another department due to immediate team needs or performance goals? G.1 Never G.2 Sometimes G.3 Often G.4 Very often
@@
Age 30-39 0.17 0.22
Age 40-49 0.40 0.34
Age >=>=> =50 0.42 0.43
Tenure <=<=< = 2 0.03 0.02
Tenure 3-5 0.06 0.04
Tenure 6-9 0.11 0.16
Tenure >=>=> =10 0.75 0.78

### v3

@@ Appendix C Theoretical Appendix
This prediction implies that greater levels of talent hoarding reduce the number of workers who apply for a promotion.
@@
This prediction implies that greater levels of talent hoarding change the composition of applicants, causing a lower share of workers with high productivity in the applicant pool.
@@
Appendix D Data Appendix
This appendix provides additional details on the data sources used in the analysis. It begins by describing the construction of the administrative dataset, which links personnel records, talent management data, and internal application data provided by the firm. It then outlines the implementation of the employee and manager surveys, including survey design, response rates, and benchmarking exercises to aid interpretation of the survey responses.
D.1 Administrative Data Used in the Analysis
This paper is based on a novel dataset combining personnel records and application data from a large manufacturing firm. To comply with the firm’s data protection protocols, all data processing was conducted within a secure data environment managed by the firm, where the dataset was assembled and anonymized.
The dataset integrates information from three internal data sources: (1) personnel records, (2) talent management data, and (3) job application data. These sources were linked using a unique employee identifier.
Personnel Records. The personnel records include detailed demographic and employment information. The following variables were extracted: employment status, date of birth, gender, marital and family status, parental leave status, date of entry into the firm, highest educational degree attained, citizenship, position, functional area, division, work location, contractual working hours, number of direct reports, reporting distance to the CEO, managerial autonomy, base pay, bonus payments, absenteeism, exit date (if applicable), retirement date, manager ID, and team ID. To create a quarterly panel dataset, the data were collapsed to the employee-by-quarter level, keeping the first observation of an employee in a given quarter.
Talent Management Data. From the firm’s internal talent management systems, I obtained information on individual performance and potential ratings, nominations to talent development programs, and worker assignments to trainings.
Job Application Data. The application tracking system provides data on all internal applications submitted by employees. Extracted variables include application date and status, application outcomes (e.g., rejection, interview, offer), vacancy ID, hiring manager ID, hierarchy level of the vacancy, as well as the division, functional area, and location associated with each vacancy.
The three data sources were linked using a unique employee id within the firm’s secure data environment.
@@ D.2 Employee Survey
All employees in my sample were invited to share their perspectives on the internal labor market at the firm. The survey yielded over 15,000 responses, corresponding to a response rate of approximately 50%. A summary of the survey content and key findings is presented in Section 3.3. This appendix section provides additional information on the survey’s implementation and instrument.
D.2.1 Survey Implementation
The survey was conducted using the firm’s established procedures for internal employee surveys. It was formally approved by the firm’s data protection officer and the workers’ council. Invitations to participate were distributed via email by the Human Resources department. Appendix Figure D1 shows the full text of the invitation email.
The survey was administered through the firm’s secure online platform. Upon accessing the survey link, respondents were prompted to provide informed consent for their responses to be recorded and analyzed. The median completion time was 13 minutes. To ensure data reliability, I restrict the analysis to respondents who spent at least five minutes completing the survey. Appendix Table D1 compares respondents to the overall population and finds that both groups are very similar in terms of their characteristics.
Figure D1: Invitation to Participate in Employee Survey
Notes: This figure presents the wording of the survey invitation sent to employees.
@@ D.2.1 Survey Implementation
D.2.2 Survey Instrument
A. Please rate following six statements. “Actively applying for positions at [Company] …” {I strongly agree, I agree, Undecided, I do not agree, I totally do not agree}
A.1 “… increases future promotion chances.”
A.2 “… does not matter since jobs are only posted proforma.”
A.3 “… would cause negative consequences by my current supervisor.”
A.4 “… is seen as disloyal to my current team.”
A.5 “… is appropriate once employees are unsatisfied with their job.”
A.6 “… should only be done after checking in with one’s direct supervisor.”
B. Which job characteristics are most important to you? Please select the two most important characteristics from the following list. {Potential for training, Potential for promotion, Pay, Flexible hours, Location, Meaningful tasks, Familiar tasks, Challenging tasks, Good relationship with colleagues, Good relationship with supervisor}
C. Did someone recommend an open position to you in the past 12 months?
C.1 Yes
C.2 No
D. You indicated that someone recommended one or more open positions to you. Who recommended these positions to you? Please select all that apply.
D.1 Colleagues within your team
D.2 Colleagues outside of your team
D.3 Your supervisor
D.4 The supervisor responsible for the open position
D.5 Your HR responsible
D.6 Family or friends
E. At the end of this survey, we are interested in your personal opinion about current challenges and potential improvements with respect to careers at [Company].
E.1 What were the reasons why you decided in the past not to apply for internal job openings at [Company]? {free-text response}
E.2 What are the main challenges that you have encountered in your career development at [Company]? {free-text response}
E.3 What are some of the ways that [Company] could be helpful to you as you are planning your career? {free-text response}
@@ D.3 Manager Survey
To provide direct evidence on managers’ incentives and talent hoarding behavior, I conducted a survey among managers in my sample. The survey achieved a response rate of 62%, resulting in over 3,000 completed responses. A summary of the survey design and key findings is provided in Section 3.4. This appendix section offers additional details on the survey implementation and questionnaire. It also presents the raw response distributions for the main survey questions used in the analysis and includes several benchmarking exercises to aid interpretation.
D.3.1 Survey Implementation
The survey was implemented using the firm’s standard procedures for internal employee surveys. Prior to launch, it received formal approval from both the firm’s data protection officer and the workers’ council. Managers were invited to participate via e-mail sent by the Human Resources department. Appendix Figure D2 displays the wording of the invitation email.
The survey was administered through the firm’s secure online survey platform. Upon accessing the survey link, respondents were prompted to provide informed consent for their responses to be recorded and analyzed. The median completion time was 14 minutes. To ensure data quality, the analysis sample is restricted to respondents who spent at least five minutes completing the survey. Appendix Table D2 shows that respondents are similar to the overall population in terms of their characteristics.
Several features suggest that the managers’ interest and trust in the survey was high. While 62% of managers responded to the survey overall, 61% responded before a reminder was sent. Appendix Table D3 compares the demographics of managers who responded before a reminder was sent (Column 1) to those who responded afterwards (Column 2). I find that the characteristics of both groups of managers are very similar. In unreported results, I also find that the likelihood to respond before the reminder was sent is not correlated with managers’ propensity to report talent hoarding using either of my two survey-based measures.
Figure D2: Invitation to Participate in Manager Survey
Notes: This figure presents the wording of the survey invitation sent to managers.
@@ D.3.1 Survey Implementation
Age 30-39 years 0.17 0.22
Age 40-49 years 0.40 0.34
Age >=>=> =50 years 0.42 0.43
Tenure <=<=< = 2 years 0.03 0.02
Tenure 3-5 years 0.06 0.04
Tenure 6-9 years 0.11 0.16
Tenure >=>=> =10 years 0.75 0.78